OpenStudy (anonymous):

complete the statement 0.3km per second=_____meters per minute my answer was 300 meter per minute

OpenStudy (anonymous):

is it correct

OpenStudy (anonymous):

i multipied by 100 so then i got 300\

OpenStudy (anonymous):

i mean 1000

hero (hero):

\[ \frac{0.3 km}{sec} = \frac{x m}{min}\] \[ \frac{0.3 (1000)m}{sec} = \frac{x m}{60 sec}\] \[ \frac{ (300)m}{sec} = \frac{x m}{60 sec}\] \[ \frac{ (300)m}{1} = \frac{x m}{60 }\] \[ \frac{ (300)}{1} = \frac{x }{60 }\] \[300(60) = x\] 18,000 = x

OpenStudy (anonymous):

i think your thinking of a different thing because thats not what where suppose to do

hero (hero):

Nope, it's the correct answer. And that's how you do it.

hero (hero):

seconds and meters both cancel with this set up. Kilo = 1000

OpenStudy (anonymous):

i dont understand is there another/ easyer way

hero (hero):

Trust me you can understand these steps

hero (hero):

Obviously, you weren't understanding the other way. If you were, you wouldn't be here.

OpenStudy (anonymous):

i was just making sure if my answer was correct i understud the other way

hero (hero):

But your answer wasn't correct.

hero (hero):

I'll post a detailed solution for you so that you understand my steps.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

oh wait i get it i get it thanks a whole lot!!!!!!

hero (hero):

Actually, I'll just explain what I did. k = kilo = 1000 1 min = 60 sec x = the blank we're trying to find the value of per means "divided by" Insert all of that into the formula I wrote above Then multiply both sides by sec to get rid of it Divide both sides by meters to get rid of it. Then cross multiply to get the value for x I hope you understand.

OpenStudy (anonymous):

yes thanks!!!

hero (hero):

I knew you could understand. Good job